OpenStudy (anonymous):

F(x)=5x+12. Find f^-1(x)

OpenStudy (anonymous):

you want the derivative?

OpenStudy (anonymous):

let y = 5x + 12 then x = (y - 12)/5 this is exactly what the inverse of f is f^-1(x) = (x - 12)/5